Solve the equation of the Linear Homogeneous Differential Equations with Constant 

Coefficients.

1. Solve y’’ – y’ – 2y = 0

2. Solve y’’ + 4y = 0

3. (D3 + 3D2 – 4D)y = 0

4. (D2 – 6D + 9)y = 0


1
Expert's answer
2021-05-21T07:31:07-0400

1) "y''-y'-2y=0"


Auxiliary equation is-


"m^2-m-2=0\\m^2-2m+m-2=0\\m(m-2)+1(m-2)=0\\(m-2)(m+1)=0\\m=2,-1"


Solution is-

"y=c_1e^{2x}+c_2e^{-x}"


"2)y''+4y=0"


Auxiliary equation is-

"m^2+4=0\\(m+2i)(m-2i)=0\\m=\pm2i"


Solution is-

"y=c_1 cos2x+c_2sin2x"


"(3)(D^3+3D^2-4D)y=0"


Auxiliary equation is-

"m^3+3m^2-4m=0\\m(m^2+3m-4)=0\\m(m^2+4m-m-4)=0\\m(m+4)(m-1)=0\\m=0,1,-4"


Solution is-

"y=c_1+c_2e^x+c_3e^{-4x}"


"(4) (D^2 \u2013 6D + 9)y = 0"


Auxiliary equation-

"m^2-6m+9=0\\(m-3)^2=0\\m=3,3"


Solution is-

"y=(c_1+c_2x)e^{3x}"